OpenStudy (anonymous):

(3/5)x-(1/10)x=x-(7/2)

OpenStudy (turingtest):

step 1) to get rid of the fractions multiply everything by 10 show me what you get after that

OpenStudy (anonymous):

why by 10 tho?

OpenStudy (anonymous):

x=7

OpenStudy (turingtest):

because it is the greatest common denominator between 5,10, and 2

OpenStudy (anonymous):

do the same as we did in previous prob

OpenStudy (anonymous):

im soo confused :(

OpenStudy (turingtest):

i.e. all of the denominators can be factored 2=2 5=5 10=2*5 so 10 will wipe out all the fractions, try it and see!

OpenStudy (anonymous):

i dont understand fractions!

OpenStudy (turingtest):

can you multiply (1/2) by 10?

OpenStudy (anonymous):

do i put it like this? (1/2) * (10/1)?

OpenStudy (turingtest):

if you want, sure so what does that simplify to?

OpenStudy (anonymous):

5?

OpenStudy (turingtest):

\[{1\over2}\times{10\over1}={10\over2}=5\]yes! now do that to all the fractions, what do you have?

OpenStudy (anonymous):

okay to 3/5 * 10 = 6 1/10*10/1=1

OpenStudy (anonymous):

*so

OpenStudy (anonymous):

7/2*10/1=35/1??

OpenStudy (anonymous):

did i do it right?

OpenStudy (turingtest):

good so far but don't bother writing the "over one" that is unnecessary, all numbers are over one, so 7/2*10/1=35 so what do you have now after simplifying\[10*{3\over5}x-10*{1\over10}x=10x-10*{7\over2}\]?

OpenStudy (turingtest):

yes you have done right so far :)

OpenStudy (anonymous):

ok, what do i do next?

OpenStudy (turingtest):

let's keep in mind what we are doing: we multiplied both sides of (3/5)x-(1/10)x=x-(7/2) by 10 and got 10[(3/5)x-(1/10)x]=10[x-(7/2)] which after simplifying the fractions gives 6x-x=10x-35 good so far?

OpenStudy (anonymous):

yesss

OpenStudy (turingtest):

good :) so now we simplify by "combining like terms" (i.e. combine x's with x's, and pure numbers with pure numbers) the first thing we can simplify is the left side 6x-x=? you tell me...

OpenStudy (anonymous):

is it just 6?

OpenStudy (turingtest):

no, combining like terms works like this: 3x+5x=(3+5)x=8x 12x-3x=(12-3)x=9x so try again, what is 6x-x=?

OpenStudy (anonymous):

but the other X has no number with it?

OpenStudy (anonymous):

unless its 5x?

OpenStudy (turingtest):

Yes! exactly! all numbers can be thought of as having an invisible one multiplying it because 1x=x so now we have 5x=10x-35 the next step is to get all the x's on one side and all the numbers without x's on the other side of the equals sign, any idea how to do that?

OpenStudy (anonymous):

do i need to subrtracts 5x-5x and 5x-10x?

OpenStudy (turingtest):

close.... you need to subtract 5x from both sides which gives 5x-5x=10x-5x-35 (you wrote above 5x-10x, so that is wrong because it is backwards) so what does the above simplify to?

OpenStudy (anonymous):

5x-35?

OpenStudy (anonymous):

is the answer 7? lol

OpenStudy (turingtest):

yes! good job. just be sure to write out the steps properly though, you need to write BOTH SIDES of the equation 0=5x-35 35=5x x=7 congrads =)

OpenStudy (anonymous):

YAY!!! hahah >:)